A sample of N2 gas (2.0 mmol) effused through a pinhole in 5.5 s. It will take __________ s for the same amount of CH4 to effuse under the same conditions.



1
Expert's answer
2022-03-25T00:56:01-0400

t directly proportional to "\sqrt{M}"

(5.5/x) = ("\sqrt{28/16}" )

(5.5/x) = 1.32

x= 5.5/1.32 = 4.17s